# The peculiar hard state behaviour of the black hole X-ray binary Swift J1727.81613

### v1

@@ Abstract
Mata Sánchez et al. (2024, 2025) presented optical observations of Swift J1727, measuring its radial velocity and dynamically confirming the presence of a black hole (). Furthermore, these authors estimated a distance of kpc using an empirical relation between binary period ( of 10.8 h) and absolute -band magnitude, together with pre-outburst photometry and optical extinction estimates. Burridge et al. (2025) reanalysed these data and supplemented them with near-UV spectra, reporting an updated distance of kpc. However, most recently, Burridge et al. (private communication) have revised this estimate to kpc, incorporating updated properties of the companion star. Our analysis adopts the most recent inferred distance range of 1.5–4.3,kpc, corresponding to the 68 equal-tailed interval. The estimate reported by Mata Sánchez et al. (2025) lies well within this range. We note that distances outside this interval would only make the source appear more extreme, and therefore do not contradict our conclusions.
In this paper, we utilise the comprehensive radio light curves presented in Hughes et al. (submitted), alongside quasi-simultaneous X-ray observations, to investigate the evolution of Swift J1727 and the connection between its accretion flow and compact jets. In Section 2, we describe the radio and X-ray data selection and analysis. Section 3 presents the track, and in Section 4 we discuss the physical interpretations.
@@ 2.1 Radio Observations
Swift J1727 was the focus of multiple extensive radio campaigns, which were compiled and presented in Hughes et al. (submitted). In this work, we use a subset of those data, obtained with MeerKAT, the Karl G. Jansky Very Large Array (VLA), and the Australia Telescope Compact Array (ATCA), as these instruments captured both the rise and decay hard states.
The relation pertains specifically to compact jet emission during the hard state. Accordingly, during the rising hard state, we include only those observations taken before the onset of radio flaring (which began on 2023 September 19; MJD 60206), after which observations with the Very Long Baseline Array revealed that the flares coincided with transient jet ejections (Wood et al., 2025). Hughes et al. (submitted) presented the entire radio light curve of Swift J1727, providing the integrated flux densities, which—during most of the decay hard state—included both unresolved emission from the compact jet (reappearing at the ‘core’ position) and spatially resolved jet ejecta (see Figure 4). The angular extent of (detectable) emission regions from the compact jet is (Wood et al., 2024), rendering it unresolved at the angular resolution of Meerkat, the VLA, and ATCA (); we measure the compact jet’s flux density by fitting only the core emission.
However, throughout the soft state, steep-spectrum (and fading) radio emission was detected at the core position. Since compact jet emission is quenched by 3 orders of magnitude in the soft state and, when detected in the hard state, exhibits a flat radio spectrum, this steep-spectrum component most likely originates from long-lived transient jet(s) that remain unresolved from the radio core (similar to those seen in MAXI J1535571; Russell et al., 2019b). As a result, following the re-brightening of the compact jet during the decay hard state, the core emission is a superposition of compact jet and ejecta emission. To account for ejecta contributions, we modelled the fading steep-spectrum component with an exponential decay222A power law model was also fit to the decay; however, given the magnitude of the measurement errors, its fit was indistinguishable from the exponential.. Extrapolating this decay to the hard state epochs—and accounting for spectral indices—we subtracted the modelled value from the observed core flux density, adopting the resulting value as the true compact jet flux density (see Appendix B).
@@ 2.3 Luminosity Conversion
Many of the radio epochs include simultaneous multi-frequency coverage. When available, we selected the flux density closest in frequency to 5 GHz. To account for spectral evolution, we calculate inter-band spectral indices from the multi-frequency data. However, most observations during the decay phase were single-frequency. To avoid systematic biases associated with intra-band spectral indices—such as the artificial spectral ‘flattening’ seen in VLA detections (e.g. Heywood et al., 2016)—we interpolate the 1.28 GHz MeerKAT data to match the ATCA and VLA observation times. This approach allows us to estimate inter-band spectral indices from the interpolated flux densities. For the ATCA observations, which were recorded simultaneously at 5.5 and 9 GHz, we verified the accuracy of the interpolated spectral index by comparing it with the ATCA-only inter-band spectral index and found good agreement (see Figure 2, second panel, star markers). We then used the ensemble of inter-band spectral indices to scale all radio flux densities to 5 GHz, adopting the spectral index measurement closest in time to each corresponding flux density. We include the final values in Table 1.
@@ 3 Results
In Figure 1, we present the evolution of Swift J1727. Due to uncertainties in its distance, we plot three curves assuming distances of 1.5, 2.6, and 4.3 kpc. For context, we include an archival sample (grey circles; Bahramian & Rushton, 2022), highlighting the archetypal radio-quiet source H174332 (blue squares; Coriat et al., 2011), as well as MAXI J1348638—which underwent the most radio-quiet evolution of all archival sources to date (green diamonds; Carotenuto et al., 2021b, 2022).
@@
accounting for uncertainties in both and using the odr function from the python package scipy.
As our monitoring spans both the rise and decay hard states, we fit the data under two scenarios (adopting kpc): (i) using both outburst phases, and (ii) using only the decay data, which captures both the steep and shallow tracks (see inset of Figure 1). In both cases, the shallow track is robustly constrained and in strong agreement yielding for (i) and for (ii), which is unsurprising, given that shallow track occurs only during the decaying hard state. The transition luminosities are similarly consistent within uncertainties, with erg s-1 for (i), and erg s-1 for (ii). The power law indices are agnostic of source distance whereas the luminosities—both at the transition to the shallow track () and when rejoining the standard track ()—scale with distance as erg s-1and erg s-1, respectively.
@@
In contrast, the inverse softhard transition tends to occur more narrowly around (Maccarone, 2003; Vahdat Motlagh et al., 2019). For Swift J1727, the X-ray luminosity at the time of this transition is ,erg s-1, implying (adopting and ). While the nominal distance yields an Eddington fraction consistent with population studies (e.g., the lognormal distribution with and from Vahdat Motlagh et al., 2019), smaller distances quickly push the transition luminosity below the observed range for BH LMXBs. This discrepancy is more significant than that introduced by the aforementioned bright hard state, thereby favouring distances at or above the nominal estimate (though we note the existence of outliers with extremely low transitional luminosities, e.g., Tomsick et al., 2014; Chauhan et al., 2019).
@@ 3.2 Radio and X-ray spectral evolution
To our knowledge, Swift J1727 is also the radio-quiet system with the best-constrained X-ray spectral evolution during its evolution. Like other BH LMXBs (both radio-quiet and radio-loud), the X-ray spectrum hardens to during the initial decay, before softening as the source fades into quiescence (top two panels of Figure 2). Archival BH LMXBs exhibit a plateau at as they approach and enter quiescence (e.g., Tomsick et al., 2001, 2004; Kong et al., 2002; Corbel et al., 2006; Plotkin et al., 2013; Reynolds et al., 2014). Given that our final epoch with spectral constraints yielded , we expect that Swift J1727 continued to soften after it became too faint for reliable spectral fitting. This places a lower limit of 16 days on the X-ray softening timescale (measured from the point of minimum ), consistent with other BH LMXBs. However, we note that the few systems with comparable constraints exhibit a broad range of softening timescales (3–90 days) (Tomsick et al., 2001; Kalemci et al., 2005; Armas Padilla et al., 2013; Plotkin et al., 2017a; Beri et al., 2019; Shaw et al., 2021). In contrast, a feature that appears unique to Swift J1727 is the apparent coincidence between the minimum and the transition onto the shallow track, suggesting a physical change in the accretion flow.
@@ 4.1 Jet-based origins
Several mechanisms have been proposed to explain the observed radio-quiet and radio-loud populations. Some attribute these differences to variations in the jet properties, which may be influenced by both intrinsic factors and observational biases. Espinasse & Fender (2018) investigated the correlation between a system’s radio spectral index and its classification as radio-loud or radio-quiet. The authors found statistically significant differences between the two populations: radio-quiet sources tended to have steeper spectral indices on average () compared to the radio-loud population (). However, applying this result to Swift J1727 is not straightforward, as its spectral index varied significantly during its radio-quiet evolution ().
@@ 4.2 Accretion-based origins
Although for Swift J1727, , , , and are not currently known—nor is any potential temporal evolution of these properties—independent measurements of any subset of these parameters can help constrain the others. For example, adopting and keV at the lower-bound distance kpc predicts (conservatively fixing ), which is inconsistent with the lowest possible mass for a stellar-mass black hole. Multiwavelength constraints on key parameters—particularly distance—are likely essential for determining the nature of this system. More broadly, such constraints across a larger sample are crucial for understanding the physical origins behind the dramatic range of transitional luminosities (i.e. spanning more than six orders of magnitude; Islam & Zdziarski, 2018; Koljonen & Russell, 2019; Carotenuto et al., 2021b).
@@ 5 Conclusion
If Swift J1727.81613 lies near the upper end of the considered distance range ( kpc), its peak 1–10 keV X-ray luminosity reaches erg s-1. Applying a typical bolometric correction factor (–), this implies a bolometric luminosity exceeding the Eddington luminosity of a black hole. In this scenario, Swift J1727 would represent the most X-ray luminous hard-state BH LMXB observed to date in the plane.
@@ Acknowledgements
AKH thanks UKRI for support. AJT acknowledges that this research was undertaken thanks to funding from the Canada Research Chairs Program and the support of the Natural Sciences and Engineering Research Council of Canada (NSERC; funding reference number RGPIN–2024–04458). FC acknowledges support from the Royal Society through the Newton International Fellowship programme (NIF/R1/211296). TDR is an INAF IAF fellow. RS acknowledges the INAF grant number 1.05.23.04.04. JvdE acknowledges a Warwick Astrophysics prize post-doctoral fellowship made possible thanks to a generous philanthropic donation, and was supported by funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101148693 (MeerSHOCKS) for part of this work. CMW acknowledges financial support from the Forrest Research Foundation Scholarship, the Jean-Pierre Macquart Scholarship, and the Australian Government Research Training Program Scholarship. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352, PI: M. Linares). RF thanks UKRI, The ERC and The Hintze Family Charitable Foundation for support. RMP acknowledges support from NASA under award No. 80NSSC23M0104. DMR and PS are supported by Tamkeen under the NYU Abu Dhabi Research Institute grant CASS.
@@ Appendix A Light Curves and Interpolation
Figure 3 presents a multi-wavelength photometric and spectral view of the 2023–2024 outburst of Swift J1727. From top to bottom: (i) the 2–20 keV flux from the Monitor of All sky X-ray Image Gas Slit Camera (MAXI/GSC, top panel; Matsuoka et al., 2009; Mihara et al., 2011); (ii) the X-ray hardness ratio, defined as the ratio of flux from the 2–6 keV and 6–20 keV MAXI/GSC fluxes (second panel); (iii) the Meerkat, VLA, and ATCA light curves originally presented in Hughes et al. (Submitted); and (iv) the inter-band radio spectral indices from simultaneous multi-frequency VLA and ATCA observations (bottom panel).
Shaded grey indicates the rise and decay hard states. The vertical dot-dashed and dashed lines mark the reported hardsoft and softhard transitions, respectively (Bollemeijer et al., 2023a, b; Podgorny et al., 2024; Russell et al., 2024). As discussed in the main text, we truncate the rise hard state earlier than the reported transition to avoid contamination from flaring activity that launches multiple jet ejecta (Wood et al., 2025). These ejections, unresolved at the angular resolution of our instruments60190.166, would be spatially coincident with the compact jet. Their presence is reflected in the radio spectral index evolution (bottom panel), which displays significant variability between the rise hard state grey region and the reported transition, indicative of a rapidly evolving population of jet components.
The radio light curves shown in Hughes et al. (Submitted) represent the total integrated flux densities. Consequently, for epochs with multiple spatial components (e.g., the compact jet and discrete ejecta), the fluxes in Figure 3 reflect the sum of multiple components. While the (included) rising hard state only includes compact jet emission, the decay hard state features bright, resolved southern-moving ejecta (Figure 4), which become the dominant radio source at late times. This leads to the plateau seen in the radio decay phase of Figure 3. For a clearer view of the compact core evolution, we present core-only light curves in Figures 5 and 6. Furthermore, we show both the measured and interpolated X-ray flux values, confirming that our linear interpolation method neither introduces nor smooths over real X-ray variability.

### v2

@@ Abstract
Mata Sánchez et al. (2024, 2025) presented optical observations of Swift J1727, measuring its radial velocity and dynamically confirming the presence of a black hole (). Furthermore, these authors estimated a distance of kpc using an empirical relation between binary period ( of 10.8 h) and absolute -band magnitude, together with pre-outburst photometry and optical extinction estimates. Burridge et al. (2025) reanalysed these data and supplemented them with near-UV spectra, reporting an updated distance of kpc. However, most recently, Burridge et al. (private communication) have revised this estimate to kpc, incorporating updated properties of the companion star. Our analysis adopts the most recent inferred distance range of 1.5–4.3 kpc, corresponding to the 68 equal-tailed interval. The estimate reported by Mata Sánchez et al. (2025) lies well within this range. We note that distances outside this interval would only make the source appear more extreme, and therefore do not contradict our conclusions.
In this paper, we utilise the comprehensive radio light curves presented in Hughes et al. (2025), alongside quasi-simultaneous X-ray observations, to investigate the evolution of Swift J1727 and the connection between its accretion flow and compact jets. In Section 2, we describe the radio and X-ray data selection and analysis. Section 3 presents the track, and in Section 4 we discuss the physical interpretations.
@@ 2.1 Radio Observations
Swift J1727 was the focus of multiple extensive radio campaigns, which were compiled and presented in Hughes et al. (2025). In this work, we use a subset of those data, obtained with MeerKAT, the Karl G. Jansky Very Large Array (VLA), and the Australia Telescope Compact Array (ATCA), as these instruments captured both the rise and decay hard states.
The relation pertains specifically to compact jet emission during the hard state. Accordingly, during the rising hard state, we include only those observations taken before the onset of radio flaring (which began on 2023 September 19; MJD 60206), after which observations with the Very Long Baseline Array revealed that the flares coincided with transient jet ejections (Wood et al., 2025). Hughes et al. (2025) presented the entire radio light curve of Swift J1727, providing the integrated flux densities, which—during most of the decay hard state—included both unresolved emission from the compact jet (reappearing at the ‘core’ position) and spatially resolved jet ejecta (see Figure 4). The angular extent of (detectable) emission regions from the compact jet is (Wood et al., 2024), rendering it unresolved at the angular resolution of Meerkat, the VLA, and ATCA (); we measure the compact jet’s flux density by fitting only the core emission.
However, throughout the soft state, steep-spectrum (and fading) radio emission was detected at the core position. Since compact jet emission is quenched by 3 orders of magnitude in the soft state (as discussed in Section 1) and, when detected in the hard state, exhibits a flat radio spectrum, this steep-spectrum component most likely originates from long-lived transient jet(s) that remain unresolved from the radio core (similar to those seen in MAXI J1535571; Russell et al., 2019b). As a result, following the re-brightening of the compact jet during the decay hard state, the core emission is a superposition of compact jet and ejecta emission. To account for ejecta contributions, we modelled the fading steep-spectrum component with an exponential decay222A power law model was also fit to the decay; however, given the magnitude of the measurement errors, its fit was indistinguishable from the exponential.. Extrapolating this decay to the hard state epochs—and accounting for spectral indices—we subtracted the modelled value from the observed core flux density, adopting the resulting value as the true compact jet flux density (see Appendix B for details). Depending on the epoch, the subtracted component constituted 25–40%, 5–12%, and 10–37% of the total for MeerKAT, ATCA, and the VLA, respectively.
@@ 2.3 Luminosity Conversion
For epochs with simultaneous multi-frequency coverage, we selected the flux density measurement closest to 5 GHz. To account for spectral evolution, we calculate inter-band spectral indices from the multi-frequency data. However, most observations during the decay phase were single-frequency. To avoid systematic biases associated with intra-band spectral indices—such as the artificial spectral ‘flattening’ seen in VLA detections (e.g. Heywood et al., 2016)—we interpolate the 1.28 GHz MeerKAT data to match the ATCA and VLA observation times. This approach allows us to estimate inter-band spectral indices from the interpolated flux densities. For the ATCA observations, which were recorded simultaneously at 5.5 and 9 GHz, we verified the accuracy of the interpolated spectral index by comparing it with the ATCA-only inter-band spectral index and found good agreement (see Figure 2, second panel, star markers). We then used the ensemble of inter-band spectral indices to scale all radio flux densities to 5 GHz, adopting the spectral index measurement closest in time to each corresponding flux density. We include the final values in Table 1.
@@ 3 Results
In Figure 1, we present the evolution of Swift J1727. Due to uncertainties in its distance, we plot three sets of points assuming distances of 1.5, 2.6, and 4.3 kpc. For context, we include an archival sample (grey circles; Bahramian & Rushton, 2022), highlighting the archetypal radio-quiet source H174332 (purple squares; Coriat et al., 2011), as well as MAXI J1348638—which underwent the most radio-quiet evolution of all archival sources to date (green diamonds; Carotenuto et al., 2021b, 2022).
@@
We account for uncertainties in both and using orthogonal distance regression, implemented via the odr function from the Python package scipy.
As our monitoring spans both the rise and decay hard states, we fit the data under two scenarios (adopting kpc): (i) using both outburst phases, and (ii) using only the decay data, which captures both the steep and shallow tracks (see inset of Figure 1). In both cases, the shallow track is robustly constrained and in strong agreement yielding for (i) and for (ii), which is unsurprising, given that shallow track occurs only during the decaying hard state. The transition luminosities are similarly consistent within uncertainties, with erg s-1 for (i), and erg s-1 for (ii). The power law indices are independent of the assumed source distance, whereas the luminosities—both at the transition to the shallow track () and when rejoining the standard track ()—scale with distance as erg s-1and erg s-1, respectively.
@@
Conversely, the inverse softhard transition tends to occur more narrowly around (Maccarone, 2003; Vahdat Motlagh et al., 2019). For Swift J1727, the X-ray luminosity at the time of this transition is erg s-1, implying (adopting and ). While the nominal distance yields an Eddington fraction consistent with population studies (e.g., the lognormal distribution with and from Vahdat Motlagh et al., 2019), smaller distances quickly push the transition luminosity below the observed range for BH LMXBs. This discrepancy is more significant than that introduced by the aforementioned bright hard state, thereby favouring distances at or above the nominal estimate (though we note the existence of outliers with extremely low transitional luminosities, e.g., Tomsick et al., 2014; Chauhan et al., 2019).
@@ 3.2 Radio and X-ray spectral evolution
To our knowledge, Swift J1727 is also the radio-quiet system with the best-constrained X-ray spectral evolution during its evolution. Like other BH LMXBs (both radio-quiet and radio-loud), the X-ray spectrum hardens to during the initial decay, before softening as the source fades into quiescence (top two panels of Figure 2). Archival BH LMXBs exhibit a plateau at as they approach and enter quiescence (e.g., Tomsick et al., 2001, 2004; Kong et al., 2002; Corbel et al., 2006; Plotkin et al., 2013; Reynolds et al., 2014). Given that our final epoch with spectral constraints yielded , we expect that Swift J1727 continued to soften after it became too faint for reliable spectral fitting. This places a lower limit of 16 days on the X-ray softening timescale (measured from the point of minimum ), consistent with other BH LMXBs. However, we note that the few systems with comparable constraints exhibit a broad range of softening timescales (3–90 days; Tomsick et al., 2001; Kalemci et al., 2005; Armas Padilla et al., 2013; Plotkin et al., 2017a; Beri et al., 2019; Shaw et al., 2021).
In contrast, a novel result from the Swift J1727 outburst is the temporal coincidence between the minimum and the onset of the shallow track. This alignment is illustrated in Figure 2a: between 2024 Apr 3–23 (MJD 60403–60423; 20 days), the X-ray flux declined by a factor of 2 as the spectrum evolved toward its minimum . Immediately thereafter, over the following 20 days (2024 Apr 23–May 14; MJD 60423–60444), the X-ray flux dropped by nearly three orders of magnitude, marking the transition onto the shallow track. To highlight this further, we use the interpolated X-ray fluxes to estimate the time range corresponding to the transitional luminosity (), based on the decay-only fit; this range is indicated by the vertical grey region in Figure 2a. A clear temporal alignment is seen between and the epoch of minimum (vertical dotted line), reinforcing the link between the spectral minimum and the onset of the shallow track. This suggests that, in the case of Swift J1727, the steep-to-shallow transition marks a genuine physical change in the accretion flow, consistent with multi-outburst analyses of H174332 (Cao et al., 2014), but here observed within a single outburst. Other sources exhibiting spectral evolution during outburst decay have not shown such transitions. For instance, MAXI J1820070 remained on the standard track throughout its evolution (Shaw et al., 2021).
@@ 4.1 Jet-based origins
Several mechanisms have been proposed to explain the observed radio-quiet and radio-loud populations. Some attribute these differences to variations in the jet properties, which may be influenced by both intrinsic factors and observational biases. Espinasse & Fender (2018) investigated the correlation between a system’s radio spectral index and its classification as radio-loud or radio-quiet. The authors found statistically significant differences between the two populations: radio-quiet sources tended to have steeper spectral indices on average () compared to the radio-loud population (). Applying this interpretation to Swift J1727 is not straightforward, as its spectral index varied significantly during the decay (), reaching optically thin values at later epochs. As discussed in Section 3.2, we cannot rule out the possibility that the late-time optically thin spectral index resulted from unsubtracted ejecta emission. However, given that Swift J1727 exhibited a flat spectral index during the early stages of the decay, if the steepening is indeed due to ejecta, this may imply that the compact jet remained self-absorbed (and spectrally flat) throughout, making Swift J1727 an outlier with respect to the findings of (Espinasse & Fender, 2018).
@@ 4.2 Accretion-based origins
Although , , , and are not currently known for Swift J1727—nor is any potential temporal evolution of these properties—independent measurements of any subset can help constrain the others. For example, adopting and keV at a lower-bound distance of kpc predicts (assuming ), which is inconsistent with even the lowest plausible mass for a stellar-mass black hole. Other combinations yield more realistic limits; for instance, and keV at kpc predict . Multiwavelength constraints on key parameters—particularly distance—are therefore essential to determine the nature of this system. More broadly, such constraints across a larger sample are critical for understanding the physical origins of the dramatic range in transitional luminosities (i.e. spanning more than six orders of magnitude; Islam & Zdziarski, 2018; Koljonen & Russell, 2019; Carotenuto et al., 2021b).
@@ 5 Conclusion
If Swift J1727.81613 lies near at or beyond the upper end of the distance range ( kpc), its peak 1–10 keV X-ray luminosity reaches erg s-1. Applying a typical bolometric correction factor (–), this implies a bolometric luminosity exceeding the Eddington luminosity of a black hole. In this scenario, Swift J1727 would represent the most X-ray luminous hard-state BH LMXB observed to date in the plane.
@@ Acknowledgements
AKH thanks UKRI for support. AJT acknowledges that this research was undertaken thanks to funding from the Canada Research Chairs Program and the support of the Natural Sciences and Engineering Research Council of Canada (NSERC; funding reference number RGPIN–2024–04458). FC acknowledges support from the Royal Society through the Newton International Fellowship programme (NIF/R1/211296). TDR is an INAF IAF fellow. RS acknowledges the INAF grant number 1.05.23.04.04. JvdE acknowledges a Warwick Astrophysics prize post-doctoral fellowship made possible thanks to a generous philanthropic donation, and was supported by funding from the European Union’s Horizon Europe research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101148693 (MeerSHOCKS) for part of this work. CMW acknowledges financial support from the Forrest Research Foundation Scholarship, the Jean-Pierre Macquart Scholarship, and the Australian Government Research Training Program Scholarship. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352, PI: M. Linares). RF thanks UKRI, The ERC and The Hintze Family Charitable Foundation for support. RMP acknowledges support from NASA under award No. 80NSSC23M0104. DMR and PS are supported by Tamkeen under the NYU Abu Dhabi Research Institute grant CASS. GRS is supported by NSERC Discovery Grant RGPIN-2021-0400.
@@ Appendix A Light Curves and Interpolation
Figure 3 presents a multi-wavelength photometric and spectral view of the 2023–2024 outburst of Swift J1727. From top to bottom: (i) the 2–20 keV flux from the Monitor of All sky X-ray Image Gas Slit Camera (MAXI/GSC, top panel; Matsuoka et al., 2009; Mihara et al., 2011); (ii) the X-ray hardness ratio, defined as the ratio of flux from the 2–6 keV and 6–20 keV MAXI/GSC fluxes (second panel); (iii) the Meerkat, VLA, and ATCA light curves; and (iv) the inter-band radio spectral indices from simultaneous multi-frequency VLA and ATCA observations (bottom panel).
Shaded grey indicates the rise and decay hard states. The vertical dot-dashed and dashed lines mark the reported hardsoft and softhard transitions, respectively (Bollemeijer et al., 2023a, b; Podgorny et al., 2024; Russell et al., 2024). As discussed in the main text, we truncate the rise hard state earlier than the reported transition to avoid contamination from flaring activity that launches multiple jet ejecta (Wood et al., 2025). These ejections, unresolved at the angular resolution of our instruments, would be spatially coincident with the compact jet. Their presence is reflected in the radio spectral index evolution (bottom panel), which displays significant variability between the rise hard state grey region and the reported transition, indicative of a rapidly evolving population of jet components.
The radio light curves shown in Hughes et al. (2025) represent the total integrated flux densities. Consequently, for epochs with multiple spatial components (e.g., the compact jet and discrete ejecta), the fluxes in Figure 3 reflect the sum of multiple components. While the (included) rising hard state only includes compact jet emission, the decay hard state features bright, resolved southern-moving ejecta (Figure 4), which become the dominant radio source at late times. This leads to the plateau seen in the radio decay phase of Figure 3. For a clearer view of the compact core evolution, we present core-only light curves in Figures 5 and 6. Furthermore, we show both the measured and interpolated X-ray flux values, confirming that our linear interpolation method neither introduces nor smooths over real X-ray variability.